Davis Construction company was a low bidder on a construction project to build and earthen dam for $2,000,000. The project was begun in 2017. Cost and other are presented bellow.

2017 2018

Costs incurred during the year $450,000 $1,100,000

Total costs 1,050,000 50,000

Percentage of completion 30%   96.875%

Billings during the year 500,000 1,400,000

Cash collections during the year 400,000 1,400,000

Assume the Beavis recognizes revenue on the contract over time according to percentage of completion

Required

1. Compute the amounts of revenue and gross profit (or loss) to be recognized in 2017 and 2018. Please label whether it is a profit or loss.
Answer #1

Answers

  • Requirement 1

2017

2018

Revenue

Revenue

$           600,000.00

Gross Profits (Loss)

Gross Profits (Loss)

$           150,000.00

  • Working

Working

Year end 2017

Year end 2018

A

Contract Price

$       2,000,000.00

$       2,000,000.00

B

Cost Incurred to Date

$           450,000.00

$       1,550,000.00

C

Estimated cost yet to be incurred to complete the contract

$       1,050,000.00

$             50,000.00

D = B+C

Total Cost

$       1,500,000.00

$       1,600,000.00

E = (B/D) x 100

% of Completion

30.00%

96.88%

F = A x E

Revenue to date

$           600,000.00

$       1,937,500.00

G

Revenue of Previous year

$                            -  

$          600,000.00

H = F - G

Net Revenue this year

$           600,000.00

$       1,337,500.00

I (=B)

Cost to date

$           450,000.00

$       1,550,000.00

J

Cost to date of previous year

$                            -  

$          450,000.00

K = I - J

Net Cost for the year

$           450,000.00

$       1,100,000.00

L = H - K

Gross Profits

$           150,000.00

$          237,500.00